OpenStudy (anonymous):

simplify the equation 6x/ 9(x+y) 3x*2/3*3(x+y) is where im stuck

jimthompson5910 (jim_thompson5910):

What term do you see cancelling?

OpenStudy (anonymous):

the 3s is that it

jimthompson5910 (jim_thompson5910):

So the answer is \[\Large \frac{2x}{3(x+y)}\] after you erase those '3's that cancel out

jimthompson5910 (jim_thompson5910):

this is because x*2 is the same as 2*x or 2x

OpenStudy (anonymous):

so it would be 2x/3x+3y

jimthompson5910 (jim_thompson5910):

either \[\Large \frac{2x}{3(x+y)}\] or \[\Large \frac{2x}{3x+3y}\]

jimthompson5910 (jim_thompson5910):

they're both equivalent since that 3 is just being distributed through

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

np